[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Aug 06, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Aug 6 10:05:56 UTC 2016


HSU Shift Log for Sat Aug 06 2016 06:05:49 GMT-0400 (EDT)

Shift: A

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: The EPAC ATCF was making a noise at the end of the previous  
shift that could have been a hard drive issue. Nothing abnormal was heard  
on this shift.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160806/c68ac34d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list